OpenStudy (anonymous):

What are the x- and y-intercepts of the line? 2x – 3y = –12 A. x-intercept 4, y-intercept –6 B. x-intercept –6, y-intercept 4 C. x-intercept –4, y-intercept 6 D. x-intercept 6, y-intercept –4

OpenStudy (♪chibiterasu):

Well first to find the x-intercept, we can replace the 3 with 0, which leaves 2x = -12 No we can divide 2 on both sides and we get -6, so the x-intercept is -6

OpenStudy (shamil98):

First put the equation in point-slope form.

OpenStudy (♪chibiterasu):

Yeah you can do that too.

OpenStudy (♪chibiterasu):

Now we do the same with y: 0x - 3y = -12 -3y = -12 Now notice that we have -3 instead of 3. This is because 3y was being subtracted from 2x. If it were being added, we would have 3y. Now divide -3 on both sides: -3y/-3 = -12/-3 y = 4 So the x intercept is -6 and the y-intercept is 4. The answer is B.

OpenStudy (anonymous):

Thnx